Three prizes are to be distributed in a Creative Design Talent Search Contest. The value of the second prize is five-sixths the value of the first prize, and the value of the third prize is fourth-fifths that of the second prize. Express the total value of the three prizes as an algebraic expression.

Answer by stanbon(75887)

The value of the second prize is five-sixths the value of the first prize, and the value of the third prize is fourth-fifths that of the second prize.
Express the total value of the three prizes as an algebraic expression.
-----------------------------------------------
Let the value of the 1st prize be "x"
Value of the 2nd prize is (5/6)x
Value of the 3rd prize is (4/5)(5/6)x = (2/3)x
---------------------------------------
Add to get:
x + (5/6)x + (2/3)x
= (13/6)x
= 2 1/6 x
===============
Cheers,
Stan H.